OpenStudy (anonymous):

Determine the slope between the points (-3, 0) and (0, 5) A. 3/5 B. 5/3 C. -5/3 D. -3/5

OpenStudy (nubeer):

slope = y2-y1/x2-x1

OpenStudy (anonymous):

it is 5/3

OpenStudy (anonymous):

slope = y2-y1/x2-x1= 5-0/0-(-3)= 5/3 that is the slope